Can you help me with this question?
Write three quadratic equations, with a,b,and c(coefficients of x^2, x, and the constant) as 1.Integers 2. Rational numbers 3. Irrational numbers

Answer by stanbon(75887)

Write three quadratic equations, with a,b,and c(coefficients of x^2, x, and the constant) as
1.Integers --------------: y = x^2 + 2x + 3
2. Rational numbers -----: y = (1/2)x^2 + (2/3)x + (1/5)
3. Irrational numbers----: y = sqrt(2)x^2 + sqrt(3)x + sqrt(7)
------------------------------------------------------------------
Cheers,
Stan H.